OpenStudy (lovelyharmonics):

I have another question c: How many grams of nitric acid will react completely with a block of iron metal that is 4.5 cm by 3.0 cm by 3.5 cm, if the density of iron is 7.87 g/mL? Show all steps of your calculation as well as the final answer. Fe + HNO3 → Fe(NO3)3 + H2 @ganeshie8

OpenStudy (lovelyharmonics):

Yay (\^-^/) chemistry!

OpenStudy (lovelyharmonics):

Are you having fun trying to learn it yet?

ganeshie8 (ganeshie8):

`4.5 cm by 3.0 cm by 3.5 cm` start by finding the total mass of the iron block

ganeshie8 (ganeshie8):

Volume of block = ?

ganeshie8 (ganeshie8):

Mass = Volume x density = ?

OpenStudy (lovelyharmonics):

Do I just multiply them together?

ganeshie8 (ganeshie8):

yess

OpenStudy (lovelyharmonics):

47.25

ganeshie8 (ganeshie8):

yes 47.25 cm^3 us the volume. what about mass ?

OpenStudy (lovelyharmonics):

So 47.25/7.87=6.0

ganeshie8 (ganeshie8):

you need to multiply : Mass = Volume * density

OpenStudy (lovelyharmonics):

Ohh that's a big difference .-. 371.8575

ganeshie8 (ganeshie8):

Yes! so we have a 371.8575 g of iron metal ready for reaction with nitric acid

ganeshie8 (ganeshie8):

lets see how much nitric acid it takes

ganeshie8 (ganeshie8):

convert 371.8575 g to moles

ganeshie8 (ganeshie8):

http://www.chemicool.com/elements/iron.html 1 mol of Fe weighs 55.85 g so how many mols are present in 371.8575 g of Fe ?

OpenStudy (lovelyharmonics):

Okay so 371.8575/6.02x10^23

ganeshie8 (ganeshie8):

nope, number of mols present in 371.8575 g of Fe = 371.8575 / 55.85

ganeshie8 (ganeshie8):

= 6.66 mol

ganeshie8 (ganeshie8):

that means we need 6.66 mol of nitric acid for the reaction as well

ganeshie8 (ganeshie8):

Find the molar weight of HNO3, then multiply by 6.66

ganeshie8 (ganeshie8):

H : 1 N : 14 O3 : 16*3 = 48 ---------------- HNO3 : 1+14+48 = 63 g

ganeshie8 (ganeshie8):

1 mol of HNO3 weight 63g

ganeshie8 (ganeshie8):

6.66 mol would weigh how much ?

OpenStudy (lovelyharmonics):

You said you don't know much about chem, you know enough to make me have a headache XD and it's 419.46

ganeshie8 (ganeshie8):

thats it ! lol this is not chemistry, this is just math.

OpenStudy (lovelyharmonics):

.-. chem is basically all math... Ad I suck at math.... That's why I had to retake both precancerous honors and chemistry honors .-. Because I can't understand them

OpenStudy (lovelyharmonics):

XD oh my god it autocorrected precac to precancerous

ganeshie8 (ganeshie8):

haha idk what autocorrect has in mind lol... you're good in math too ! but im not seeing you much math recently hmm